# MINDS: Detection of an inner gas disk caused by evaporating bodies around HD 172555

### v2

@@ Appendix H Estimate of Stellar UV and X-ray Fluxes
The potential X-ray emission from HD 172555 can be constrained using observations from Chandra. Schröder et al. (2008) report a non-detection, placing an upper limit on the X-ray luminosity of erg s-1, assuming a coronal temperature of 0.9 keV ( K), a value rather high for an A-type star. For a cooler corona with K, more typical of A-type stars, the upper limit becomes erg s-1, as simulated using XSPEC (Arnaud, 1996) with an APEC coronal model and solar abundances, applying the Chandra ACIS-I response over the standard 0.1–2.4 keV energy band.
@@
We conclude that erg s-1 (0.1–10 keV) is a likely estimate for the X-ray luminosity of Pic. Although caution is warranted due to potential UV contamination, the Chandra detection at a similar flux level supports this result. By analogy, this value can also serve as a plausible upper limit for HD 172555, given the Chandra non-detection and the relatively insensitive upper limits available unless unrealistically high coronal temperatures are assumed. This inferred luminosity is consistent with the published Chandra upper limit of erg s-1 for HD 172555.

### v3

@@ Appendix H Estimate of Stellar UV and X-Ray Fluxes
The potential X-ray emission from HD 172555 can be constrained using observations from Chandra. Schröder et al. (2008) reported a nondetection, placing an upper limit on the X-ray luminosity of erg s-1, assuming a coronal temperature of 0.9 keV ( K), a value rather high for an A-type star. For a cooler corona with K, more typical of A-type stars, the upper limit becomes erg s-1, as simulated using XSPEC (Arnaud, 1996) with an APEC coronal model and solar abundances, applying the Chandra ACIS-I response over the standard 0.1–2.4 keV energy band.
@@
We conclude that erg s-1 (0.1–10 keV) is a likely estimate for the X-ray luminosity of Pic. Although caution is warranted due to potential UV contamination, the Chandra detection at a similar flux level supports this result. By analogy, this value can also serve as a plausible upper limit for HD 172555, given the Chandra nondetection and the relatively insensitive upper limits available unless unrealistically high coronal temperatures are assumed. This inferred luminosity is consistent with the published Chandra upper limit of erg s-1 for HD 172555.